Consider the differential operator T:P3(R)→P3(R) given by

T(p(x))=2p″(x)−2p′(x)−2p‴(x)

Find an ordered basis F for P3(R) such that T acts like a shift operator with respect to F, i.e. M??(?)(1 point) Consider the differential operator T : P3(R) → P3(R) given by T(p(x)) = 2p(x) — 2p(x) – 2p(x) 1 Find an ordered
Answer #1

an²+ bx?text de B = -bax² + (129-4bx + 4b-20-1291 then t cant bit (xtd) = 2 [asit byt (xt d) & 2 [ax& box & Coxtd.) - 2 canㅋ aco, b20, GB -20 120 = 0 벌 Hence 300 2 (;A=0b) 22 - 4 New Tpa(11) = P310 ㅋ -67 + [2a - 4b)xt Ub-2C-12a = ub-212020. 12a -Ub

answered by: ANURANJAN SARSAM